# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v2

@@ Abstract
In this paper, we study the Lagrangian aspects of the inhomogeneous incompressible Euler equation (IIE in short). We establish a geodesic description of this equation and discuss the associated geometric structures. We also find the derivation of IIE from the Hamilton-Pontryagin action principle and derive the corresponding Lagrangian formulation. Appealing to this Lagrangian perspective, we prove Lagrangian analyticity of IIE.
@@ 1 Introduction
The geometry of hydrodynamics has been extensively studied since the seminal work of Arnold [Ar66] and Ebin-Marsden [EM70], where the motion of perfect fluids in a smooth bounded domain $D$ is interpreted as geodesic in the infinite dimensional Riemannian manifold $\mathsf{SDiff}(D)$ of volume-preserving diffeomorphisms of $D$, which is equipped with the right-invariant Riemannian metric given by the kinetic energy of the fluid. This geometric perspective has been extensively generalized since then; we refer to [AK98][HMR98][MR99] for further development of this topic, and [KMM20] for a recent survey on geometric hydrodynamics.
From a mechanical viewpoint, Arnold’s geodesic formulation is a natural infinite-dimensional generalization of the classical least action principle in Lagrangian mechanics. Flipping to the Hamiltonian side, we have an analogous Hamilton-Pontryagin type variational principle for hydrodynamic models (see [BCHM00][Ho21]), which leads to the Lagrangian-Eulerian formulation of fluid equations [Co01a]. We remark that the above perspective admits a natural generalization which effectively applies to viscous hydrodynamic models. We refer to [CI08] for the Navier-Stokes case and to [MP25] for an extended discussion. However, all the above discussions are restricted to fluid models with mechanical Lagrangians, where kinetic and potential energy are separated.
A natural example which lies out of the zoo of mechanical Lagrangian is the following inhomogeneous version of the incompressible Euler equation (IIE in short), which is an extensively studied model in past decades:
@@
The above equation describes fluids with variable density, transported by its own incompressible velocity field. The study of the above equation was initiated in the celebrated monograph [[Lio96]], see also [Da06][[Da10]][[DF11]]. The main focus of the works listed above is to address well-posedness issues in various functional settings, analyzing IIE from an Eulerian viewpoint using delicate Harmonic analysis tools. Contrary to the well-known Yudovich theorem of 22-$d$ Euler’s equation, the global well-posedness issue of IIE in 22d is still open. There are also works devoted to studying the Beale-Kato-Majda (cf.[[BKM84]] )type blow up/continuation criteria for IIE in the past few years, we refer the readers to [[BLS20]][[Fa25]].
@@
However, different from the analytical side, the geometric structure and Lagrangian perspective of the inhomogeneous Euler equation is much less studied. To the author’s knowledge, the only relevant literature in this direction is [[Mar76]], which addressed the geodesic formulation of IIE on $\mathsf{SDiff}(D)$, as a generalization of [EM70]. To fill the gap in the literature, in this paper we establish a Lagrangian approach to the analysis of IIE, which stems from its geometric structure. We discuss the Hamiltonian formulation of the equation and derive a Lagrangian-Eulerian representation of the solution of IIE. The advantages of such Lagrangian perspective are two-folds: First, the Lagrangian formulation motivates us to find a vorticity formulation of IIE, which totally eliminates the pressure; Second, benefited from the Lagrangian perspective, we obtain an explicit recursive formula for Taylor coefficients (in time) of Lagrangian displacement, and accordingly, demonstrate Lagrangian analyticity of the solution.
@@
The structure of this paper is organized as follows: In section 2, we study the geodesic formulation of IIE with respect to a weighted Riemannian metric on $\mathsf{SDiff}(D)$, where we interpret the weighted Leray projector as the second-fundamental form. The sectional curvature and Lagrangian stability are also discussed. In section 3, we derive IIE from a Hamilton-Pontryagin type variational principle, and obtain a Lagrangian-Eulerian formulation of IIE as a Lagrangian formulae for solution of the Hamiltonian ODE given by the variational principle. We also derive the vorticity formulation from our geometric construction, and recover the local existence results in our framework. Finally in section 4, using our Lagrangian representation, we prove the analyticity of Lagrangian trajectories of the IIE.
In this paper, we will use the letter $C$ to denote generic positive constants which may change from line to line. $A\lesssim B$ means there exists $C>0$ such that $A\leq CB$.
@@ Acknowledgements
The author would like to thank V.Vicol for stimulating discussions which initiate the study of Lagrangian analyticity of IIE. The author thanks T.Drivas for pointing out the ongoing work [DG26+] and for helpful discussions. The author thanks B.Khesin, A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. This work was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453 (PI Mazzucato).
@@ 2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of $\mathsf{SDiff}(D)$ which naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished notes [DrWS25], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension (The $\mathsf{SDiff}(D)$ setting).
@@
We start with the well-known Jacobi-Liouville formula which will be repeatedly used in this section.
Lemma 2.1
@@
Let $A:[0,1]\to\mathsf{GL}(d)$ be a smooth curve in the general linear group $\mathsf{GL}(d)$, then we have:
$\frac{d}{dt}\det(A_{t})=\mathsf{Tr}(A^{-1}\dot{A})\det(A)$ (2)
$\frac{d}{dt}A_{t}^{-1}=-A_{t}^{-1}\dot{A}_{t}A_{t}^{-1}$ (3)
Proof. To show (2), let $\{M_{ij}\}$ be minors of $A=\{a_{ij}\}$, we expand for each fixed $i$:
$\det(A)=\sum_{j=1}^{d}(-1)^{i+j}M_{ij}a_{ij}$
Using the fact that $M_{ij}$ is independent of $a_{ij}$, we have
$\frac{\partial\det(A)}{\partial a_{ij}}=(-1)^{i+j}M_{ij},\quad\frac{d}{dt}\det(A)=\sum_{i,j=1}^{d}\frac{\partial\det(A)}{\partial a_{ij}}\frac{da_{ij}}{dt}=\sum_{i,j=1}^{d}(-1)^{i+j}M_{ij}\frac{da_{ij}}{dt}$
Since we have
$(A^{-1})_{ij}=\frac{1}{\det(A)}(-1)^{i+j}M_{ij}$
We conclude
$\frac{d}{dt}\det(A)=\det(A)\sum_{i,j=1}^{d}(A^{-1})_{ij}\frac{da_{ij}}{dt}=\mathsf{Tr}(A^{-1}\dot{A})\det(A)$
(3) is a simple consequence of the chain rule:
$0=\frac{d}{dt}(A_{t}A_{t}^{-1})=\dot{A}_{t}A_{t}^{-1}+A_{t}\frac{d}{dt}A_{t}^{-1}$
Rearranging the above, we obtain (3). ∎
Remark 2.2
As pointed out by [Kh25], yet another simpler proof of the above lemma is by Abel’s theorem on Wronskian of matrix ODEs. More precisely, $\det A$ is the Wronskian of matrix $A$, which satisfies $\dot{A}=AA^{-1}\dot{A}$, then apply Abel’s theorem to coefficient matrix $\dot{A}A^{-1}$ the desired result follows.
@@ 2.1 The Riemannian Structure
Let $D=\mathbb{T}^{d}$, which is the physical domain of our fluids, equipped with natural periodic boundary conditions. Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$ ($\mathfrak{X}_{\sigma}(D)$) the space of (divergence-free) smooth vector fields on $D$. We remark that the formulation below also works in the setting of smooth compact Riemannian manifold, and works under $H^{s+1}$ Sobolev regularity with $s>d/2$, after suitable technical modification. Here, we only consider the smooth case on torus for simplicity.
Our geometric description relies on the interpretation of $\mathsf{SDiff}(D)$ as a level set submanifold of the full diffeomorphism group $\mathsf{Diff}(D)$ with infinite dimension and co-dimension, and the inhomogeneous Euler equation will be derived as geodesic on the submanifold, with respect to some weighted metric. This formalism closely follows [DrWS25]. (See also the ongoing work [DG26+], dealing with dynamics on rapidly oscillating constraint submanifold in infinite dimensions.)
More precisely, we reformulate
@@
$\nabla\cdot(\delta X\circ X^{-1})](X(a))=0.$
@@
$T_{X}\mathsf{SDiff}(D)=\{v\circ X:\nabla\cdot v=0\}.$
Now we move on to define the Riemannian metric on $\mathsf{SDiff}(D)$. We denote by $\mathsf{Dens}(D)$ the space of smooth densities in $D$. Fix a reference density $\rho_{0}\in\mathsf{Dens}(D)$ which is required to avoid concentration or vacuum, that is, satisfying the following assumption in this paper:
Assumption (A):
There exists $C_{1}>1$ such that
$0<C_{1}^{-1}\leq\rho_{0}(a)\leq C_{1}<\infty\quad\text{ for all }a\in D.$
@@
here $\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A$. Clearly $\rho$ is a rearrangement of $\rho_{0}$, so the assumption (A) is preserved under the action of $\mathsf{SDiff}(D)$ on $\mathsf{Dens}(D)$.
@@
$\mathcal{N}_{X}:\varphi\in H^{1}(D)\to\rho_{0}^{-1}\nabla\varphi(X)=\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(\varphi\circ X)\in N_{X}\mathsf{SDiff}(D).$ (6)
@@
$\mathcal{N}_{X}^{*}(\mu):\delta X\in T_{X}\mathsf{Diff}(D)\to\langle\mu,\nabla\cdot(\delta X\circ A)\rangle_{*}.$ (7)
which satisfies the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{X}^{*}(\mu))$.
Moreover, in this context, the musical isomorphism (inertia operator in the Language of mechanics) is $\alpha\to\rho_{0}^{-1}\alpha$, since for any $1$-form $\alpha$ we have
$\langle\delta X,\alpha\rangle_{*}=\mathsf{G}_{X}(\alpha_{\sharp},\delta X)=\langle\delta X,\alpha_{\sharp}\rangle_{L_{\rho_{0}}^{2}}.$
Which implies that $\alpha_{\sharp}=\rho^{-1}\alpha$. Hence, the vector representative $\mathcal{N}_{X}$ of covector $\mathcal{N}_{X}^{*}$ is:
$\mathcal{N}_{X}(\mu):=\rho_{0}^{-1}\mathcal{N}_{X}^{*}(\mu)=-\rho_{0}^{-1}\nabla\mu(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(X_{*}\mu).$
which should be interpreted in the distributional sense.
The following proposition is a consequence of standard computation:
Proposition 2.3
@@
    $(Jf)_{a,a^{\prime}}:=\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$ (10)
@@
    $(Jf)_{a,a^{\prime}}^{-1}(X)=G_{\rho}(X(a),X(a^{\prime})).$ (11)
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$, under mean-zero condition.
Proof. The proof follows from straightforward computation. For (I) we have
@@
$(Jf)_{a,a^{\prime}}^{-1}(X)=-G_{\rho}(X(a),X(a^{\prime})),$
where $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$.

### v3

@@ Abstract
In this paper, we study the inhomogeneous incompressible Euler equation (IIE in short) from a Lagrangian perspective. We establish a geodesic description of this equation and discuss the associated geometric structures. We also find the derivation of IIE from the Hamilton-Pontryagin action principle and derive the corresponding Lagrangian formulation. A byproduct is a new vorticity formulation of IIE. We also prove the Lagrangian analyticity of IIE using our Lagrangian representation formula.
@@ 1 Introduction
The geometry of hydrodynamics has been a vivid research direction since the seminal work of Arnold [2] and Ebin-Marsden [18], where the motion of perfect fluids in a smooth bounded domain $D$ is interpreted as geodesic in the infinite dimensional Riemannian manifold $\mathsf{SDiff}(D)$ of volume-preserving diffeomorphisms of $D$, which is equipped with the right-invariant Riemannian metric given by the kinetic energy of the fluid. This geometric perspective has been extensively generalized since then; we refer to [1][24][30] for further development of this topic, and [26] for a recent survey on geometric hydrodynamics including compressible fluids.
From a mechanical viewpoint, Arnold’s geodesic formulation is a natural infinite-dimensional generalization of the classical least action principle in Lagrangian mechanics. Flipping to the Hamiltonian side, we have an analogous Hamilton-Pontryagin type variational principle for hydrodynamic models (see [6][25]), which leads to the Lagrangian-Eulerian formulation of fluid equations [12]. We remark that the above perspective admits a natural generalization which effectively applies to viscous hydrodynamic models. We refer to [10] for the Navier-Stokes case and to [33] for an extended discussion. However, all the above discussions are restricted to fluid models with mechanical Lagrangians, where velocity and position dependence are decoupled and appear only in kinetic and potential energy respectively.
A natural example which falls outside the class of mechanical Lagrangians is the following inhomogeneous version of the incompressible Euler equation (IIE in short), which is an extensively studied model in past decades:
@@
The above equation describes the motion of incompressible fluids with variable density. Here, the inhomogeneity of the fluid forces the kinetic energy density to depend on both position and velocity variables:
$\frac{1}{2}\lvert u\rvert^{2}\rho=\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X^{-1},\quad\dot{X}=u(X).$
@@
The above equation was systematically studied in the celebrated monograph [28]. Analysis of the equation has been done in various functional settings in a series of recent works, see [14][15][13]. The main focus of the works listed above is to address the well-posedness issues in various function spaces and to analyze IIE from an Eulerian viewpoint using Harmonic analysis tools. However, contrary to the well-known Yudovich theorem of 22-$d$ homogeneous Euler’s equation, the global well-posedness issue of IIE in 22d is still open. In recent years, there have also been works devoted to Beale–Kato–Majda-type (cf.[5] ) blow-up/continuation criteria for IIE, we refer the readers to [3][9][21].
@@
However, different from the analytical side, the geometric structure and Lagrangian perspective of the inhomogeneous Euler equation is much less studied. To the best of the author’s knowledge, the only relevant literature in this direction is [32], which studies the geodesic formulation of IIE on $\mathsf{SDiff}(D)$, as a generalization of the celebrated geodesic formulation of the incompressible Euler equation. The Hamiltonian aspects and Lagrangian formulation (a Weber-type formula, see [12]) of IIE, to the best of the author’s knowledge, are absent in the current literature.
To fill the gap in the literature, in this paper we establish a Lagrangian approach to the analysis of IIE, which stems from geometric consideration. We discuss the Hamiltonian formulation of the equation and derive a Lagrangian-Eulerian representation of the solution of IIE. The advantages of such a Lagrangian perspective are twofold: First, the Lagrangian formulation motivates us to find a vorticity formulation of IIE, which totally eliminates the pressure; Second, benefiting from the Lagrangian perspective, we obtain an explicit recursive formula for the time-series coefficients of the Lagrangian displacement, and demonstrate Lagrangian analyticity of the solution accordingly.
We believe that the vorticity formulation we obtained for IIE may uncover new approaches to the study of IIE. The key takeaway here is that: Geometrically, vorticity in incompressible fluid models is naturally given by the curl of the momentum rather than the velocity, due to the dual Lie transport (coadjoint motion) nature of momentum, together with the following commutation relation, which encodes the vorticity transport:
$\nabla\times\mathscr{L}_{u}^{*}=\mathscr{L}_{u}\nabla\times,\quad\mathscr{L}_{u}^{*}v:=u\cdot\nabla v+\nabla^{*}uv.$
Thus, there is a natural difference between the vorticity of homogeneous and inhomogeneous quantities, based on geometric considerations. As the vorticity formulation plays an important role in the study of various problems of incompressible Euler equation, for instance the finite time blow up, continuation criteria and stability of steady states, the author believes that our vorticity formulation may be helpful in investigating analogous problems for IIE. In our companion paper [35], based on our vorticity formulation, we study the lifespan and continuation criteria of IIE via a Lagrangian approach.
@@
The structure of this paper is organized as follows: In section 2, we study the geodesic formulation of IIE with respect to a weighted Riemannian metric on $\mathsf{SDiff}(D)$, where we interpret the weighted Leray projector as the second-fundamental form. The sectional curvature and Lagrangian stability are also discussed. In section 3, we derive IIE from a Hamilton-Pontryagin type variational principle, and obtain a Lagrangian-Eulerian formulation of IIE as a Lagrangian formula for the solution of the Hamiltonian ODE given by the variational principle. We also derive the vorticity formulation from our geometric construction, and recover the local existence results in our framework. Finally, in section 4, using our Lagrangian representation, we prove the time-analyticity of Lagrangian trajectories of the IIE.
@@ Acknowledgements
The author thanks V.Vicol for stimulating discussions which initiate the study of Lagrangian analyticity of IIE. The author thanks T.Drivas for pointing out the ongoing work [16] and for helpful discussions. The author thanks B.Khesin, H.N.Lopes, M.Lopes, A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. This work was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453.
@@ 2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of the volume-preserving diffeomorphism group, which is equipped with a weighted Riemannian metric and naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished notes [17], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension.
2.1 Some Conventions
We fix some frequently used notations. Let $D$ be the physical domain of our fluids. Unless otherwise specified, we assume $D=\mathbb{T}^{d}$, that is the unit box $0,1]^{d}$ equipped with periodic boundary condition. We remark that most our results also hold in the setting of smooth compact Riemannian manifold or $\mathbb{R}^{d}$, and works under $H^{s+1}$ Sobolev regularity with $s>d/2$, after suitable technical modification (e.g. replacing mean-zero condition with slip boundary condition for compact manifold $D$, or with decay condition for $D=\mathbb{R}^{d}$).
Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$, $\mathfrak{X}_{0}(D)$, $\mathfrak{X}_{\sigma}(D)$ the space of smooth vector fields, mean-free vector fields and divergence-free vector fields on $D$.
In this section, we will focus on the geometric formulation of IIE and only consider the smooth case on torus for simplicity. Later, for analytical results like well-posedness and analyticity, we will work in Sobolev regularity class. We will use the notation $H_{0}^{k}(D)$ for space of mean-zero $H^{k}(D)$ functions, and we will frequently use the notation $H^{s}$ without specifying the domain and range of the function. We denote by $\mathbf{P}$ the standard Leray projector on $\mathfrak{X}(D)$, which projects vector fields to its div-free part. In compact form:
$\mathbf{P}v=v-\nabla\Delta^{-1}\nabla\cdot v.$
It is well-known that $\mathbf{P}$ is a zero-order multiplier that preserves the mean value, Hence, it is bounded on each Sobolev space $H_{0}^{k}(D)$ and $\mathbf{P}:\mathfrak{X}_{0}(D)\to\mathfrak{X}_{0}(D)\cap\mathfrak{X}_{\sigma}(D)$.
We will use the letter $X$ for elements in $\mathsf{SDiff}(D)$, and letter $A$ for the inverse of $X$, which is the back-to-label map associated with flow map $X$ in fluid mechanics context.
We denote by $\mathsf{Dens}(D)$ the space of smooth densities in $D$, that is:
$\mathsf{Dens}(D)=\bigg\{\rho\in C^{\infty}(D):\int_{D}\rho(x)dx=1\bigg\}.$
We will fix a reference density $\rho_{0}\in\mathsf{Dens}(D)$, which is required to satisfy the following assumption throughout this paper:
Assumption 2.1 (Assumption (A)).
There exists $C_{1}>1$ such that
$0<C_{1}^{-1}\leq\rho_{0}(a)\leq C_{1}<\infty\quad\text{ for all }a\in D.$
The above assumption 2.1 is posed to avoid concentration or vacuum in fluid density. We denote by
$\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A,\quad X\in\mathsf{SDiff}(D),$
which is the dynamical density transported by some rearrangement $X$. We notice that the assumption 2.1 is propagated under the action of rearrangements. Hence, also propagated by transport equation. We may sometimes add the subscript $X_{t}$ to highlight its time dependence.
We will frequently deal with variable coefficient second-order elliptic operators. Fix a $\rho\in\mathsf{Dens}(D)$, we consider the following weighted Laplacian operator and the associated Poisson equation:
$\Delta_{\rho}\varphi:=\nabla\cdot(\rho^{-1}\nabla\varphi)=f\quad\text{ in }D.$
Assumption 2.1 guaranties that $\Delta_{\rho}$ is an uniformly elliptic operator. Here, we assume the forcing term $f$ satisfies the mean-zero condition:
$\int_{D}f(x)dx=0.$
Then, $\Delta_{\rho}$ is invertible in the space of mean-zero smooth functions. We will denote by $\mathcal{G}_{\rho}$ the Green’s function associated with $\Delta_{\rho}$, and the following holds:
$\int_{D}\mathcal{G}_{\rho}(x,y)dx=0,\quad\mathcal{G}_{\rho}(x,y)=\mathcal{G}_{\rho}(y,x),\quad\varphi(x)=\int_{D}\mathcal{G}_{\rho}(x,y)f(y)dy.$
We will use the letter $C$ to denote positive constants which may depend on additional parameters and change from line to line. $A\lesssim B$ means there exists $C>0$ which might depend on other harmless parameters such that $A\leq CB$.
@@ 2.1 Some Conventions
We end this subsection with the well-known Jacobi-Liouville formula, which will be repeatedly used in the sequel.
Lemma 2.2.
@@
Let $\mathbf{M}:[0,1]\to\mathsf{GL}(d)$ be a smooth curve in the general linear group $\mathsf{GL}(d)$, then we have:
$\frac{d}{dt}\det(\mathbf{M})=\mathsf{Tr}(\mathbf{M}^{-1}\dot{\mathbf{M}})\det(\mathbf{M});$ (2)
and
$\frac{d}{dt}\mathbf{M}^{-1}=-\mathbf{M}^{-1}\dot{\mathbf{M}}\mathbf{M}^{-1}.$ (3)
Proof of Lemma 2.2.
We first notice (3) is a simple consequence of the chain rule:
$0=\frac{d}{dt}(\mathbf{M}\mathbf{M}^{-1})=\dot{\mathbf{M}}\mathbf{M}^{-1}+\mathbf{M}\frac{d}{dt}\mathbf{M}^{-1}.$
Rearranging the above equality, we obtain (3).
To show (2), the standard approach is to use minor expansion. Here, we provide a simple proof suggested by [27].
Notice that $\det\mathbf{M}$ is the Wronskian of matrix $\mathbf{M}$, since the matrix ODE that $\mathbf{M}$ satisfies reads
$\dot{\mathbf{M}}=\mathbf{M}(\mathbf{M}^{-1}\dot{\mathbf{M}})=:\mathbf{M}\tilde{\mathbf{M}}$,
we apply the Abel’s formula for Wronskian (c.f. [37]) to conclude:
$\frac{d}{dt}\det\mathbf{M}=\mathsf{Tr}(\tilde{\mathbf{M}})\det\mathbf{M}_{t}=\mathsf{Tr}(\mathbf{M}^{-1}\dot{\mathbf{M}})\det\mathbf{M}.$
@@ 2.2 The Riemannian Structure
Our geometric description relies on the interpretation of $\mathsf{SDiff}(D)$ as a level set submanifold of the full diffeomorphism group $\mathsf{Diff}(D)$ with infinite dimension and co-dimension, and the inhomogeneous Euler equation will be derived as geodesic on the submanifold, with respect to some weighted metric. This formalism closely follows [17]. (See also the ongoing work [16], dealing with dynamics on rapidly oscillating constraint submanifold in infinite dimensions.)
More precisely, we rewrite $\mathsf{SDiff}(D)$ as follows:
@@
$\nabla\cdot(\delta X\circ X^{-1})](X(a))=0,$
@@
$T_{X}\mathsf{SDiff}(D)=\{v\circ X:v\in\mathfrak{X}_{\sigma}(D)\}.$
@@
$\mathcal{N}_{X}:\varphi\in H^{1}(D)\to-\rho_{0}^{-1}\nabla\varphi(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(\varphi\circ X)\in N_{X}\mathsf{SDiff}(D).$ (6)
@@
$\mathcal{N}_{X}^{*}(\mu)\in T_{X}^{*}\mathsf{SDiff}(D),\quad\mathcal{N}_{X}^{*}(\mu):=\delta X\in T_{X}\mathsf{Diff}(D)\to\langle\mu,\nabla\cdot(\delta X\circ A)\rangle_{*}.$
which in distributional sense reads:
$\mathcal{N}_{X}^{*}(\mu)=-\nabla\mu(X).$
Moreover, the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{X}^{*}(\mu))$ is satisfied.
Moreover, in this context, the musical isomorphism (inertia operator in the Language of mechanics) is $\alpha\to\rho_{0}^{-1}\alpha$. Hence, the vector representative $\mathcal{N}_{X}$ of covector $\mathcal{N}_{X}^{*}$ is:
$\mathcal{N}_{X}(\mu):=\rho_{0}^{-1}\mathcal{N}_{X}^{*}(\mu)=-\rho_{0}^{-1}\nabla\mu(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(X_{\sharp}\mu),$
which should be interpreted in the distributional sense. Here, $X_{\sharp}\mu$ is the push-forward of measure $\mu$ by diffeomorphism $X$.
In the sequel, we will fix our notation that $A$ will always denote the spatial inverse of diffeomorphism $X$. The following proposition calculates the gradient, Hessian and Gram operator of constraint functions $f_{a}$ in this setting.
Proposition 2.3.
@@
    $(Jf)_{a,a^{\prime}}:=\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=-\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$ (9)
@@
    $(Jf)_{a,a^{\prime}}^{-1}(X)=\mathcal{G}_{\rho}(X(a),X(a^{\prime})).$ (10)
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $\mathcal{G}_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$, under mean-zero condition.
Proof of Proposition 2.3.
The proof follows from straightforward computation. For (I) we have
@@
$(Jf)_{a,a^{\prime}}^{-1}(X)=-\mathcal{G}_{\rho}(X(a),X(a^{\prime})),$
where $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$. ∎